Prev Next

SW&IT Meeting Minutes 2022-04-11

Attendees: Brandon, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220404
LR20220404 - Bec, Matthieu 40.0
LR20220404 - Mechtley, Brandon 80.0
LR20220404 - Eychaner, Glenn 40.0
LR20220404 - Edmeades, David 40.0
LR20220404 - Grenz, Paul 4.0
LR20220404 - Hooper, Stephen 0.0
LR20220404 - Zhang, Yang 40.0
LR20220404 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • lbti.org domain?

Team reports

MB

  • Last week
    • OC, Management, ATT
    • JC AO report
    • TMS gmt/lbt round-up Fri.
    • TCS/AOS/DD ncpa keyword elucidations
    • PyQt3 centos7
    • OT/PIT to web, update from Gemini (mostly rewriting all of it, ready 2024 maybe)
  • This week
    • AO/sw/system planning meeting (weekly Fri @11am)
    • ARGOS github synch (Lorenzo?)
    • aorev_git dx asm, others ctd.
    • elab, JC AO report
    • TCS release for eng nights

YZ

  • Previous:
    • GMT-LBT TMS workshop on Friday.
    • LBC bias level work with Alessandro and Olga on Sunday.
    • Some TCS/OSS/Argos testing with Paul.
    • Some issuetrak investigation and discussions.
  • Next:
    • Science archive meeting?
    • Meeting with Olga on LBC non-sidereal tracking issues.
    • Some other issues investigation.
    • Resume some MODS and TMS software work.

DE

  • Dentist this morning
  • Last week:
    • Laptop followups, all good so far
    • VPN setup/verification. Silva says all good.
    • No licencing yet
    • M1 Mac Mini spec and purchase
  • This week
    • Licensing, hopefully
    • Set up Mac if it arrives
    • Office trip re: Mac laptop(s), any other items that come up

GE

  • Last week:
    • Rewrote (again) the migration of the LUCI observer interface as a Plan of Intent.
    • Read TCS documentation on GitHub and in the Wiki, and continued poking around my TCS VM to upgrade from 2018A to the latest version.
    • Yet more investigation of changes in RMI (Activation) and XML (JAXB) support in JVMs after version 8, as it impacts the LUCI GUI PoI.
    • Started working on ITs related to the Observer Scheduler GUI (8571, 8529).
  • This week:
    • Finish upgrading my TCS VM to the latest version.
    • Put together a working group on upgrading the LUCI GUIs. This may overlap with changes to the PIT and OT.
    • Work on Observer Scheduler GUI (8571, 8529).

BM

  • Last week:
    • AO SW meeting: new test plans (interrupted offset swap pauses, idlsystem thread safety, uninterruptible offsets)
    • dxlbti:
      • fixed some dxlbti confcalib/installation issues (.L/.R, auxloops, keepup)
      • made sense of hosts for sx/dx soul, lbti, adsec (was confused about ts8/ts8dx mostly)
    • argos: went over scripts with gustavo (pam/prm, using spacetrack.org, aircraft avoidance)
    • set_z0 tests: maybe Doug didn't have FAKESLOPE enabled for the second day of tests, but did for the first day.
    • looking at code for BCU power on, disturbance enabling
  • This week:
    • IT 8602 - copleyctrl fix for error reporting rather than timeout
    • IT 8553 - soulsx bcu timeouts on setup
    • ao-supervisor 109 - msgd disconnect loading disturbances
    • dxlbti bcu2k not responding (configured as 172.16.1.20)
    • discuss interrupted offsets w/JCG, MB, et al

PSG

  • Last week
    • Vacation
    • Test ARGOS proxy, white light source
    • Reconnect feature worked correctly.
    • Small glitch, restarting proxy seemed to fix, can't replicate.
    • brief SHARK-NIR work (configurable buffer size).
  • This week
    • Schedule another test of OSS ARGOS functionality,
    • Determine plan for GCS testing (new version).
    • Continue reconfigurable image buffer size work.

SH

  • LBTI data
  • Tucson machine room, bit of a cleanup and in my office as well.
  • hypervisor Tucson, vm2 not accepting upgraded controller
  • Gave Tucson surplus a gander
  • Room 507 upgrades and cleanup
  • Tucson network issues

next:

  • VMware
  • Core switches
  • Room 507 upgrades
  • Tucson network
  • Dentist 04/13
Topic revision: r10 - 11 Apr 2022, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback